# Gland Segmentation using SAM with Cancer Grade as a Prompt

### v1

@@ 2.6 Loss Function
When fine-tuning SAM, mean square error (MSE) loss was used. The concept of weight map, originally introduced in U-Net [14], is applied to the gland annotation when training both gland and contour branch of proposed model. The weight map of gland annotation is multiplied element-wise with MSE loss for each pixel. The result is summed to obtain the weight MSE loss.
@@ 3.4 Results on Whole Slide Images
Fig. 4 shows the prompted SAM-H prediction results on the GlaS dataset’s test set. It reveals that the proposed model performed well on the test sets, and there is a substantial similarity between the model’s prediction and the provided ground truth. Glands with both regular and irregular shapes are accurately predicted, and in cases where adjacent glands are very close to each other, they are successfully separated.
@@ 4 Conclusion
As benign and malignant glands exhibit varying morphological characteristics, a deep learning model can take advantage of cancer grade information as prior knowledge. A new method to provide grade prompt is designed. Heat map highlighting regions generated by Grad-CAM++ serve as a prompt to guide the segmentation model. A prompt adapter is proposed to help model adapt to the heat map prompt. By increasing the number of prompt encoders and mask decoders to two, the model can make predictions for gland and contour segmentation at the same time. Removing the overlapping area between glands and contours can help separate adjacent glands. A stepwise training approach enables the model to achieve improved training outcomes.

### v2

@@ 2.6 Loss Function
When fine-tuning SAM, mean square error (MSE) loss was used. The concept of weight map, originally introduced in U-Net [14], is applied to the gland annotation when training both gland and contour branch of proposed model. The weight map of gland annotation is multiplied element-wise with MSE loss for each pixel. The result is summed to obtain the weighted MSE loss.
@@ 3.4 Results on Whole Slide Images
Fig. 4 shows the prompted SAM-H prediction results on the GlaS dataset’s test sets. It reveals that the proposed model performed well on the test sets, and there is a substantial similarity between the model prediction and the provided ground truth. Glands with both regular and irregular shapes are accurately predicted, and in cases where adjacent glands are very close to each other, they are successfully separated.
@@ 4 Conclusion
As benign and malignant glands exhibit varying morphological characteristics, a deep learning model can take advantage of cancer grade information as prior knowledge. A new method to provide grade prompt is designed. Heat maps highlighting regions generated by Grad-CAM++ serve as a prompt to guide the segmentation model. A prompt adapter is proposed to help model adapt to the heat map prompt. By increasing the number of prompt encoders and mask decoders to two, the model can make predictions for gland and contour segmentation at the same time. Removing the overlapping area between glands and contours can help separate adjacent glands. A stepwise training approach enables the model to achieve improved training outcomes.